Chem 162 Experiment 2 Spectrophotometric Determination of Ion Concentration by Beers Law Pre-Lab Questions a. A sample absorbs 80% of the light passed through it at a certain wavelength. What is the absorbance of the sample at this wavelength? b. What would the absorbance be at the same wavelength if the sample were diluted by half? Assume the same size sample cell was used. 2, A 0.070 M solution has an absorbance ofA = 0.689. If a 1.00 cm path length was used. calculate the molar absorptivity constant () for this solution. Be sure to include the units! 3. Use the equation for the best fit line for the plot of Absorbance vs concentration below to calculate the concentration of the standard that corresponds to an absorbance value of 0.25 Beers Law Plot 34 27x+0.0032 R-0.9982 04 01 0002 0004 0006 0008 001 0012 0014。。 4. Be sure to have all of your data tables prepared in your lab notebook. You must have your volume (Part A) and concentration (Part B) calculations completed in your lab book before you come into the lab. Reproduce those tables here and hand them in as part of your prelab
Answer #1

(1)

(a)

Using relation:

A = 2 - log(%T)

Since 80% is absorbed, so %T = 100-80 = 20

Putting values we get:

A = 2-log(20) = 0.699

(b)

Assuming that diluted by half means that sample conc was decreased to half its original value:

According to Beer Lambert's law:

A = e*l*C

So, if conc decreases by half, so the absorbance will also decrease by half.

So, new absorbance = 0.699/2 = 0.3495

Hope this helps !